OpenStudy (anonymous):

If y varies directly as x and y=9 when x=-3, find y when x=18.... Please show steps on how you did it :)

jimthompson5910 (jim_thompson5910):

y varies directly as x means that y = k*x where k is some fixed number

jimthompson5910 (jim_thompson5910):

it tells us that when x =-3, y = 9, so this means y = k*x 9 = k*(-3) k = ???

jimthompson5910 (jim_thompson5910):

what did you get for k

OpenStudy (anonymous):

k=-3

OpenStudy (anonymous):

is that right?

jimthompson5910 (jim_thompson5910):

yes it is

jimthompson5910 (jim_thompson5910):

since k = -3 and y = kx, we can say that the equation is y = -3x

jimthompson5910 (jim_thompson5910):

when x = 18, what is y?

OpenStudy (anonymous):

-54

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome